OpenStudy (1018):

how to get the parallel line of this? x=-9+6t y=6+7t z=3

OpenStudy (amistre64):

strip off the direction vector

OpenStudy (amistre64):

your coeffs of t define your direction vector

OpenStudy (anonymous):

x=-9+6t y=6+7t z=3 + 0t

OpenStudy (1018):

so is the parallel line the coefficients of t?

OpenStudy (1018):

or the other ones?

OpenStudy (amistre64):

no, the parallel part of the line is defined by the coeffs of t, we have to anchor it to a point for it to define a line

OpenStudy (amistre64):



OpenStudy (1018):

ok can you teach me what to do

OpenStudy (amistre64):

strip .... off ... the coeffs .... of t what is it you are trying to do?

OpenStudy (anonymous):

there are many parallel lines to your given line

OpenStudy (1018):

oh sorry yeah i already got the answer thanks!

OpenStudy (anonymous):

you said find *the* parallel line?

OpenStudy (1018):

the whole question was find the distance etc etc

OpenStudy (amistre64):

im going to go out on a limb and say that wasnt the whole question :) but good luck

OpenStudy (1018):

sorry for the mistake i thought it wouldnt make a difference thanks anyway i got it

OpenStudy (anonymous):

amistre , just checking, the set of parallel lines is x=a+6t y=b+7t z=c